# Robo-Instruct: Simulator-Augmented Instruction Alignment For Finetuning CodeLLMs

### v1

@@ A.3 Analysis of the Generated Datasets
We first compute and plot the distribution of token lengths in the Self-Instruct-generated dataset and the Robo-Instruct-generated dataset, as shown in Fig. 6(a). Next, we measure the cosine similarity between each dataset and the RoboEvalbenchmark tasks following the approach in Magicoder [38], as depicted in Fig. 6(b). Finally, Tab. 3 presents the n-gram diversity score of each dataset, along with the number of synthesized locations and objects. Our findings indicate that both distributions and dataset statistics are very similar, suggesting that Robo-Instruct enhances the quality of the generated data over Self-Instruct rather than merely aligning the dataset towards the benchmark tasks.
A.4 RoboEvalSeed Task Example
1# Instruction: Go to Arjun’s office,
2# ask him if he is ready to head out,
3# and come back and tell me what he said
4def task_program():
@@ A.4 RoboEvalSeed Task Example
Listing 1: Seed Task Example 1
1# Instruction: Ask Alice if she needs 1, 2, or 3 boxes.
2# Go to the storage room and ask if they have that many boxes.
3# If so, go place the boxes in Alice’s office.
4# Otherwise, tell Alice you could not get the boxes.
5def task_program():
@@
Listing 2: Seed Task Example 2
1# Instruction: Check if there is a red marker in the main
2# office, and if so, tell Eve that there is a marker there.
3# If not, go to the supply room and
4# bring a red marker to the main office.
5def task_program():
@@
Listing 3: Seed Task Example 3
1# Instruction: Check every classroom if there is a whiteboard.
2# Go to Aiden’s office to tell him which room does not
3# have a whiteboard. Come back and tell me task is completed.
4def task_program():
@@
Listing 4: Seed Task Example 4
1# Instruction: Go to the kitchen and wait for someone
2# to show up. When someone shows up, ask them to open
3# the fridge, then pick up a diet coke.
4# Finally, put the diet coke in the living room.
5def task_program():
@@
Listing 5: Seed Task Example 5
1# Instruction: Take a bed sheet from the laundry room
2# and put it in each of the bedrooms.
3def task_program():
@@
Listing 6: Seed Task Example 6
A.5 Prompts to Generate Synthetic Dataset Using Self-Instruct
You are a helpful assistant. Here is a robot that has the following capabilities:
- def get_current_location() -> str:
- def get_all_rooms() -> list[str]:
- def is_in_room(object : str) -> bool:
- def go_to(location : str) -> None:
- def ask(person : str, question : str, options: list[str]) -> str:
- def say(message : str) -> None:
- def pick(obj: str) -> None:
- def place(obj: str) -> None:
Generate an interesting robot task that can be accomplished using the above capabilities.
{{SEED EXAMPLE}}
Generate an interesting robot task that can be accomplished using the above capabilities.
Table 4: Prompts to Generate Synthetic Dataset Using Self-Instruct.
A.6 CoT Prompts for InstAlign
### Role: You are an expert at understanding robot programs. You will be given a task instruction and robot program pair. However, the instruction may not align with the program well. You need to correct the task instruction to match the given robot program.
### Context: The robot only has access to the following 8 APIs and standard Python functions
- def get_current_location() -> str:
- def get_all_rooms() -> list[str]:
- def is_in_room(object : str) -> bool:
- def go_to(location : str) -> None:
- ask(person : str, question : str, options: list[str]) -> str:
- say(message : str) -> None:
- def pick(obj: str) -> None:
- def place(obj: str) -> None:
### Inputs
Original Instruction: This is a task instruction that may not align with the robot program Robot Program: This is a python function starting with ‘def task_program():‘
### Task:
1. Write down all the provided APIs used in the program and explain the effect of each API in this program
@@ 2. Examine these APIs and write down step by step what the program does
3. Combine all the results above and rewrite the instruction under # Final Corrected Instruction: You need to be specific and clear in your final corrected instruction.
Table 5: CoT Prompts for InstAlign.

### v2

@@ A.3.3 Analysis of Generated Dataset
Since RoboSim does not rely on pre-defined simulation environments, we aim to assess the diversity of programs generated by Self-Instruct and whether Robo-Instruct can maintain this diversity. To do so, we measure the number of distinct entities, such as synthetic locations and objects. As shown in Tab. 4, with a dataset of only 5,000 samples, approximately 1,000 unique objects and locations are generated, highlighting that conventional robot simulations with pre-defined environments are insufficient. Additionally, Tab. 4 presents the n-gram diversity scores for each dataset, indicating that both distributions and dataset statistics are highly similar. This suggests that Robo-Instruct not only preserves but enhances the quality of generated data compared to Self-Instruct, rather than simply aligning the dataset with benchmark tasks.
A.4 Prompts
A.4.1 RoboEval Seed Task Example
Seed Task Example 1:
1# Instruction: Go to Arjun’s office, ask him if he is ready to head out, and come back and tell me what he said
2def task_program():
@@ A.4.1 RoboEval Seed Task Example
Seed Task Example 2:
1# Instruction: Ask Alice if she needs 1, 2, or 3 boxes. Go to the storage room and ask if they have that many boxes. If so, go place the boxes in Alice’s office. Otherwise, tell Alice you could not get the boxes.
2def task_program():
@@
Seed Task Example 3:
1# Instruction: Check if there is a red marker in the main office, and if so, tell Eve that there is a marker there. If not, go to the supply room and bring a red marker to the main office.
2def task_program():
@@
Seed Task Example 4:
1# Instruction: Check every classroom if there is a whiteboard. Go to Aiden’s office to tell him which room does not have a whiteboard. Come back and tell me task is completed.
2def task_program():
@@
Seed Task Example 5:
1# Instruction: Go to the kitchen and wait for someone to show up. When someone shows up, ask them to open the fridge, then pick up a diet coke. Finally, put the diet coke in the living room.
2def task_program():
@@
Seed Task Example 6:
1# Instruction: Take a bed sheet from the laundry room and put it in each of the bedrooms.
2def task_program():
@@
A.4.2 Prompts to Generate Synthetic Dataset Using Self-Instruct
1You are a helpful assistant. Here is a robot that has the
2following capabilities:
3- def get_current_location() -> str:
4- def get_all_rooms() -> list[str]:
5- def is_in_room(object : str) -> bool:
6- def go_to(location : str) -> None:
7- def ask(person : str, question : str, options: list[str]) -> str:
8- def say(message : str) -> None:
9- def pick(obj: str) -> None:
10- def place(obj: str) -> None:
11Generate an interesting robot task that can be accomplished using the
12above capabilities.
13{SEED EXAMPLE 1}
14
15...
16
17Generate an interesting robot task that can be accomplished using the
18above capabilities.
19{SEED EXAMPLE 6}
20
21Generate an interesting robot task that can be accomplished using the
22above capabilities.
A.4.3 CoT Prompts for InstAlign
1### Role
2You are an expert at understanding robot programs.
3You will be given a task instruction and robot program pair.
4However, the instruction may not align with the program well.
5You need to correct the task instruction to match the given robot program.
6
7### Context
8The robot only has access to the following 8 APIs and
9standard Python functions
10- def get_current_location() -> str:
11- def get_all_rooms() -> list[str]:
12- def is_in_room(object : str) -> bool:
13- def go_to(location : str) -> None:
14- ask(person : str, question : str, options: list[str]) -> str:
15- say(message : str) -> None:
16- def pick(obj: str) -> None:
17- def place(obj: str) -> None:
18
19### Inputs
20Original Instruction
21 This is a task instruction that may not align with the robot program
22Robot Program
23 This is a python function starting with def task_program():‘
24
25### Task
261. Write down all the provided APIs used in the program and
27explain the effect of each API in this program
@@ 23 This is a python function starting with def task_program():‘
293. Combine all the results above and rewrite the instruction
30You need to be specific and clear in your final corrected instruction.
A.5 Real-world Experiments
A.5.1 Program Generation Results Beyond RoboEval
We conducted a small qualitative experiment to assess the performance of the base model, Self-Instruct fine-tuned models, Robo-Instruct fine-tuned models, and GPT-4 on long-horizon tasks. To showcase performance, we created two new long-horizon tasks that differ significantly from those in the RoboEval benchmark. Using a temperature setting of 0, we generated programs with each model and found that our Robo-Instruct fine-tuned model successfully produced correct programs, while both the base model and Self-Instruct fine-tuned model failed. Additionally, GPT-4 made an error on the second instruction. We show the generation results from Robo-Instruct fine-tuned models and GPT-4 below.
Task instruction 1:
1Lets play a game: Double and give it to the next person. Start with 1 dollar. Go to rooms A, B, C, D, E, F, and G. If you see someone, tell them how much money you have. Then ask if they would like to take the money now or double the amount and give it to the next person. If they choose to take it, the game is over, and you should come back to me. Otherwise, double your money and continue. If, in the end, no one takes the money, tell me how much you still have.
Robo-Instruct fine-tuned model (Ours) [Success]
1def task_program():
2 start_loc = get_current_location()
3 list_of_rooms = ["A", "B", "C", "D", "E", "F", "G"]
4 money = 1
5 for room in list_of_rooms:
6 go_to(room)
7 if is_in_room("person"):
8 say("I have \$" + str(money))
9 response = ask("", "Would you like to take the money now
10 or double the amount and give it to the next person?",
11 ["Take now", "Double and give"])
12 if response == "Take now":
13 go_to(start_loc)
14 say("The game is over, I gave the money to " + room)
15 return
16 else:
17 money *= 2
18 go_to(start_loc)
19 say("No one took the money, I still have \$" + str(money))
GPT-4 [Success]
1def task_program():
2 start_loc = get_current_location()
3 rooms = ["A", "B", "C", "D", "E", "F", "G"]
4 money = 1
5 for room in rooms:
6 go_to(room)
7 if is_in_room("person"):
8 say("I have " + str(money) + " dollars")
9 response = ask("", "Would you like to take the money now
10 or double the amount and give it to the next person?",
11 ["Take it", "Double it"])
12 if response == "Take it":
13 say("Game over. The person in room " + room + " took the money.")
14 go_to(start_loc)
15 return
16 else:
17 money *= 2
18 go_to(start_loc)
19 say("No one took the money. I still have " + str(money) + " dollars.")
Task instruction 2:
1Go to my office and check if I have a table, a chair, and a monitor there. If any of these items are missing, go to Jasons office and see if he is there. If he is, ask him if I can borrow the missing items. If he agrees, pick up each missing item and bring it to my office. If Jason is not in his office or he says no, come back and tell me the
2reason.
Robo-Instruct fine-tuned model (Ours) [Success]
1def task_program():
2 start_loc = get_current_location()
3 go_to("my office")
4 table_found = is_in_room("table")
5 chair_found = is_in_room("chair")
6 monitor_found = is_in_room("monitor")
7 if not table_found:
8 go_to("Jason’s office")
9 if is_in_room("Jason"):
10 response = ask("Jason", "Can I borrow the table from you?",
11 ["Yes", "No"])
12 if response == "Yes":
13 go_to("Jason’s office")
14 pick("table")
15 go_to("my office")
16 place("table")
17 else:
18 go_to(start_loc)
19 say("Jason said no")
20 else:
21 go_to(start_loc)
22 say("Jason is not in his office")
23 if not chair_found:
24 go_to("Jason’s office")
25 if is_in_room("Jason"):
26 response = ask("Jason", "Can I borrow the chair from you?",
27 ["Yes", "No"])
28 if response == "Yes":
29 go_to("Jason’s office")
30 pick("chair")
31 go_to("my office")
32 place("chair")
33 else:
34 go_to(start_loc)
35 say("Jason said no")
36 else:
37 go_to(start_loc)
38 say("Jason is not in his office")
39 if not monitor_found:
40 go_to("Jason’s office")
41 if is_in_room("Jason"):
42 response = ask("Jason", "Can I borrow the monitor from you?",
43 ["Yes", "No"])
44 if response == "Yes":
45 go_to("Jason’s office")
46 pick("monitor")
47 go_to("my office")
48 place("monitor")
49 else:
50 go_to(start_loc)
51 say("Jason said no")
52 else:
53 go_to(start_loc)
54 say("Jason is not in his office")
55 go_to(start_loc)
56 say("task is completed")
GPT-4 [Failure]
1def task_program():
2 start_loc = get_current_location()
3 go_to("my office")
4 missing_items = []
5 for item in ["table", "chair", "monitor"]:
6 if not is_in_room(item):
7 missing_items.append(item)
8 if len(missing_items) > 0:
9 go_to("Jason’s office")
10 if is_in_room("Jason"):
11 for item in missing_items:
12 response = ask("Jason", "Can I borrow your " + item +
13 "?", ["Yes", "No"])
14 if response == "Yes":
15 # BUG: did not return to Jason’s office after going back to my office
16 pick(item)
17 go_to("my office")
18 place(item)
19 else:
20 go_to(start_loc)
21 say("Jason did not allow me to borrow his " + item)
22 return
23 else:
24 go_to(start_loc)
25 say("Jason is not in his office")
26 return
27 go_to(start_loc)
28 say("Task is completed")
A.5.2 Generation Latency
(a) Example Program 1 generated by our fine-tuned model.
(b) Example Program 1 generated by GPT-4.
Figure 11: Our fine-tuned model is approximately 6x faster in inference speed than GPT-4 (Part 1).
(c) Another Example Program generated by our fine-tuned model.
(d) Another Example Program generated by GPT-4.
Figure 12: Our fine-tuned model is approximately 6x faster in inference speed than GPT-4 (Part 2).
A.6 Toy Examples Beyond Service Mobile Robots
A.6.1 Robot with low-level controls
Consider a tabletop manipulation scenario with a potential API function, is_rotate(robot_gripper_ name, radians), where the robot’s gripper has a physical constraint, allowing rotation only within the range $[-\frac{\pi}{6},\frac{\pi}{6}]$ radians. For the following generated program snippet:
1def task_program():
2 rotate("left hand", math.pi/6)
3 rotate("left hand", math.pi/6)
4 rotate("left hand", math.pi/6)
DynamicEval will first infer that ”left hand” is an entity of the robot gripper type. Then, regardless of the initial configuration of the gripper, DynamicEval will throw an error because the program causes the gripper to exceed its allowable range of motion.
A.6.2 AI-powered personal digital assistant
Consider a broader application than robotics: code generation for an AI-powered personal digital assistant. This AI assistant could handle scheduling events using an API function like schedule_on_calendar(event, start_time, duration). Given the instruction: ”My schedule is free tomorrow morning. Please create two 1-hour timeslots for office hours for my robotics and deep learning class.” The assistant could generate a program to create these timeslots:
1def task_program():
2 schedule_on_calendar("robotics class office hour",
3 "9:30 am", "1 hr")
4 schedule_on_calendar("deep learning class office hour",
5 "10:00 am", "1 hr")
In this example, DynamicEval needs to reason about the entities “robotics class office hour” and “deep learning class office hour”, which are categorized as event types. The event type indicates that these entities have associated timeslots. The state of these entities is defined by the time they occur: robotics class office hour is set for 9:30-10:30 am, and deep learning class office hour is set for 10:00-11:00 am. During evaluation, DynamicEval can identify a time conflict between these two office hours and thus determine that the generated program is invalid.